The relativistic momentum of a neutron(rest mass energy=940 MeV)whose kinetic energy k=200 MeV is given by
1
Expert's answer
2019-05-13T10:47:51-0400

under the circumstances relativistic momentum 

"p=\frac{T\sqrt{1+\frac{2E_0}{T}}}{c}= 1.1\cdot10^{-19}\sqrt{1+2\cdot\frac{94}{20}}=3.5\cdot10^{-19} (\frac{kgm}{s})"